# Prompts Are Programs Too! Understanding How Developers Build Software Containing Prompts

### v1

@@ 5.1.6. Evaluation
Many participants’ tasks often involved custom tasks that did not have a clear, quantified notion of correctness; as such, evaluating the prompt was a challenging task (Parnin et al., 2023; Dolata et al., 2024). In developing ML-enabled systems, many of the evaluation metrics focus on quantitative measures (Wan et al., 2019; Giray, 2021; Nahar et al., 2022), while prompt engineering includes manual inspection (Gero et al., 2022; Mishra et al., 2023). Participants performed manual testing of the prompt’s outputs simply by qualitatively assessing the outputs (P2, P4, P6, P7, P8, P12, P13, P14, P16, P17, P18, P19, P20), since it was difficult to “quantify the concept of good” (P10) for outputs. For example, one participant who created a voice memo summarization tool did the following to test his tool: “I would just sit in the chair and talk to my phone for like 10 minutes just to fine-tune and see, ‘Okay, does it ask me the right question? Does the summary make sense?”’ (P2).
@@
Manual testing was difficult to scale and quantify progress (P6, P15): “It is extremely hard to realize the impact of the [changes] without it going through some manual intervention, where there is a person who is very finely checking out the differences” (P15), especially in the face of regressions (see Section 4.1.3). Thus, several participants supplemented this assessment with quantitative metrics via programmatic testing of the prompt for a scaled evaluation (P1, P5, P8, P9, P11, P13, P14, P16, P18): “When I update or change something, I run it manually on 3 to 4 examples… In cases where I have a proper evaluation setup, I run on 50 to 100 cases. Those 50 cases make me more confident” (P18). This often involved using other FMs to generate scores for qualitative constructs (P4, P7, P9, P10, P15, P19), such as safety (Furmakiewicz et al., 2024): “I have a scoring mechanism with…an evaluation agent [where the] score is from 0 to 100 [for groundedness]… Whenever [the agent] is failing, I can shortlist those examples and then manually go through them” (P19). However, FM-based evaluations could also introduce errors: “[The evaluation agent] itself is an LLM, right? It’s only 90% reliable. So 10% of the time, if the test case is failing, it’s because the evaluation agent wasn’t able to score it correctly” (P19).
Testing occurs at different scopes.
@@
Participants noted testing the prompts at varying scopes. For applications that involved multiple prompts, participants tested individual components of the prompt program (P8, P16, P19, P20), akin to unit testing: “Usually, each test case that we had written was aimed at testing one specific part of the prompt” (P20). Participants also described performing integration testing (P2, P8, P9, P11, P12, P13, P14, P16, P18, P19): “We evaluated each of [the prompts] separately, and now we are evaluating that as a collection” (P11). However, this type of testing was difficult, as combining prompts could yield unexpected interactions (see Section 4.1.3).
@@ 5.2. Implications
Developing ML-enabled systems requires a variety of expertise like software engineering, data science, and math (Wan et al., 2019; Giray, 2021). Yet, the expertise is captured across multiple team members, which can introduce collaboration challenges (Nahar et al., 2022). Even though prompt programming does involve both ML and programming knowledge, this expertise is now centralized in a single individual. This could reduce friction between different roles. However, in addition to traditional software engineering expertise, like debugging, testing, or knowledge in specific programming languages (Liang et al., 2022; Baltes and Diehl, 2018), prompt programmers must develop new skills as they engage in novel activities like building datasets and debugging model errors with data (Qian et al., 2024). We notice that by becoming ML practitioners, prompt programmers encounter many of the long-standing challenges contended by the ML and NLP communities, like model drift (Chen et al., 2023), language (Joshi et al., 2020) and cultural biases (Santy et al., 2023; Pei and Jurgens, 2023) in datasets and models, and annotator subjectivity (Aroyo and Welty, 2015). Close collaboration between the software engineering and ML communities is paramount to learn techniques to address these challenges in prompt programming.
@@
The rapid iteration involved in prompt programming means that prompt programmers must make sense of a deluge of information. Each iteration, the participant must track the prompt, the change being made, the FM’s output, and associated metrics. However, participants reported having trouble making sense of all of this information (P3, P6, P18, P19). Therefore, future work should investigate better ways for prompt programmers to version and navigate through prompt iterations and their associated artifacts, such as datasets and model outputs. Exploratory programming tools (e.g., Variolite (Kery et al., 2017)) could provide inspiration, as exploratory programming is also highly iterative and unsystematic in nature (Kery and Myers, 2017).
6. Discussion & Conclusion
Generative pre-trained models (e.g., GPT-4 (Achiam et al., 2023)), also known as foundation models (FMs), have allowed programmers to write natural language prompts that power intelligent AI experiences (Parnin et al., 2023) in software products. We study this phenomenon, prompt programming, by interviewing 20 developers engaged in prompt development across a variety of contexts. We use Straussian grounded theory to develop a better understanding of the prompt programming process. We identify 14 observations on prompt programming (see Section 1), which indicate prompt programming is significantly different compared to traditional software development. This produces several implications for software engineering educators, practitioners, researchers, and tool makers.
For educators
The lack of standardized best practices and unreliable mental models of prompt programmers pose significant challenges in teaching prompt programming. Participants said having prior experience with FMs and “prompt intuition” (P15) made it easier to develop and debug prompt programs. Since this knowledge is tacit, software educators could teach students with assignments that give students experience with prompt programming to develop their prompt intuition. Additionally, software engineering educators could consider teaching data science skills, such as dataset curation, data cleaning, data shaping, and debugging model errors with data (Kim et al., 2016; Qian et al., 2024).
For practitioners
Practitioners should consider identifying potential fairness issues at the start of prompt development, rather than after deployment. This could be done by developing data collection pipelines from culturally diverse participants, following Santy and Liang et al. (Santy et al., 2023). Additionally, practitioners should focus on developing benchmarks to catch performance regressions. These benchmarks should account for paraphrasings of the same query and be evolving to address training-serving skew (Nahar et al., 2022). Finally, practitioners should make collaboration a top concern due to the tight coupling of chained prompts. Consider regular check-ins with colleagues and frequent integration testing of chained prompts during development.
@@ 6. Discussion & Conclusion
“[Prompt programming] will require any traditional software developer to adjust their mindset… It is much more scientific and experimental than traditional software development, and the results are not instant” (P15). Participants struggled at each step while developing prompt programs, from requirements to evaluation, as existing developer tools were not designed to handle the challenges of prompt programming. We follow Hassan et al. (2024) and Parnin et al. (2023) in underscoring the importance of developer tools for prompt programming. Such explorations have begun in earnest (e.g., Mishra et al., 2023; Wu et al., 2022a; Jiang et al., 2022; Arawjo et al., 2023). Future tool ecosystems should support the full range of activities in prompt programming, from interfaces for rapid prototyping, dataset curation, and data sensemaking, to regression testing, prompt chaining, collaboration, and more.
@@
To facilitate replication of the study, we provide our supplemental materials (Authors, 2024), which is available on FigShare at https://figshare.com/s/de7da35d6351ce374995. This includes our survey instrument, interview protocol, codebook, and the list of papers reviewed during the focused literature review. We will make these materials publicly available upon acceptance.
@@ Acknowledgements.
We thank our participants—without whom this work would not be possible—for their wonderful insights. We also extend our gratitude to Chenyang Yang and Nadia Nahar for the feedback on the work throughout its development. Last but not least, we give a special thanks to MeiReferences
@@
- Arawjo et al. (2023) Ian Arawjo, Priyan Vaithilingam, Martin Wattenberg, and Elena Glassman. 2023. ChainForge: An open-source visual programming environment for prompt engineering. In ACM Symposium on User Interface Software and Technology (UIST). 1–3. https://doi.org/10.1145/3586182.3616660
- Aroyo and Welty (2015) Lora Aroyo and Chris Welty. 2015. Truth is a lie: Crowd truth and the seven myths of human annotation. AI Magazine 36, 1 (2015), 15–24.
- Arpteg et al. (2018) Anders Arpteg, Björn Brinne, Luka Crnkovic-Friis, and Jan Bosch. 2018. Software engineering challenges of deep learning. In Euromicro Conference on Software Engineering and Advanced Applications (SEAA). IEEE, 50–59. https://doi.org/10.1109/SEAA.2018.00018
- Authors (2024) Anonymous Authors. 2024. Supplemental Materials to ”Prompts Are Programs, Too! Towards a Theory on Prompt-Based Programming with Foundation Models”. The supplemental materials can be found on HotCRP or on Figshare: https://figshare.com/s/de7da35d6351ce374995.
- Baltes and Diehl (2018) Sebastian Baltes and Stephan Diehl. 2018. Towards a theory of software development expertise. In ACM Joint Meeting on European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE). 187–200. https://doi.org/doi/10.1145/3236024.3236061
- Bender et al. (2021) Emily M Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell. 2021. On the dangers of stochastic parrots: Can language models be too big?. In ACM conference on Fairness, Accountability, and Transparency (FAccT). 610–623.
- Betker et al. (2023) James Betker, Gabriel Goh, Li Jing, Tim Brooks, Jianfeng Wang, Linjie Li, Long Ouyang, Juntang Zhuang, Joyce Lee, Yufei Guo, et al. 2023. Improving image generation with better captions. Computer Science. https://cdn.openai.com/papers/dall-e-3. pdf 2, 3 (2023), 8.
- Brown (2020) Tom B Brown. 2020. Language models are few-shot learners. arXiv preprint arXiv:2005.14165 (2020).
@@
- Chavan et al. (2024) Omkar Sandip Chavan, Divya Dilip Hinge, Soham Sanjay Deo, Yaxuan Wang, and Mohamed Wiem Mkaouer. 2024. Analyzing developer-ChatGPT conversations for software refactoring: An exploratory study. In International Conference on Mining Software Repositories (MSR). 207–211. https://doi.org/10.1145/3643991.3645082
- Chen et al. (2023) Lingjiao Chen, Matei Zaharia, and James Zou. 2023. How is ChatGPT’s behavior changing over time? arXiv preprint arXiv:2307.09009 (2023).
@@
- Cole and Gillies (2022) Tom Cole and Marco Gillies. 2022. More than a bit of coding:(un-) Grounded (non-) theory in HCI. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI). 1–11. https://doi.org/10.1145/3491101.3516392
- Corbin and Strauss (2015) Juliet Corbin and Anselm Strauss. 2015. Basics of qualitative research. Vol. 14. Sage.
- Dijkstra (1982) Edsger W Dijkstra. 1982. On the role of scientific thought. Selected writings on computing: a personal perspective (1982), 60–66. https://doi.org/10.1007/978-1-4612-5695-3_12
- Dijkstra (2005) Edsger W Dijkstra. 2005. On the foolishness of ”natural language programming”. Program Construction: International Summer School (2005), 51–53. https://doi.org/10.1007/BFb0014656
- Dolata and Crowston (2023) Mateusz Dolata and Kevin Crowston. 2023. Making sense of AI systems development. IEEE Transactions on Software Engineering (TSE) (2023). https://doi.org/10.1109/TSE.2023.3338857
- Dolata et al. (2024) Mateusz Dolata, Norbert Lange, and Gerhard Schwabe. 2024. Development in times of hype: How freelancers explore generative AI?. In IEEE/ACM International Conference on Software Engineering (ICSE). 1–13. https://doi.org/10.1145/3597503.3639111
@@
- Giray (2021) Görkem Giray. 2021. A software engineering perspective on engineering machine learning systems: State of the art and challenges. Journal of Systems and Software 180 (2021), 111031. https://doi.org/10.1016/j.jss.2021.111031
- Glaser and Strauss (2017) Barney Glaser and Anselm Strauss. 2017. Discovery of grounded theory: Strategies for qualitative research. Routledge.
@@
- Hassan et al. (2024) Ahmed E Hassan, Dayi Lin, Gopi Krishnan Rajbahadur, Keheliya Gallaba, Filipe Roseiro Cogo, Boyuan Chen, Haoxiang Zhang, Kishanthan Thangarajah, Gustavo Oliva, Jiahuei Lin, et al. 2024. Rethinking software engineering in the era of foundation models: A curated catalogue of challenges in the development of trustworthy FMware. In Companion Proceedings ACM International Conference on the Foundations of Software Engineering. 294–305. https://doi.org/10.1145/3663529.3663849
@@
- Jacovi and Goldberg (2020) Alon Jacovi and Yoav Goldberg. 2020. Towards faithfully interpretable NLP systems: How should we define and evaluate faithfulness?. In Annual Meeting of the Association for Computational Linguistics (ACL), Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel Tetreault (Eds.). 4198–4205. https://doi.org/10.18653/v1/2020.acl-main.386
@@
- Jiang et al. (2022) Ellen Jiang, Kristen Olson, Edwin Toh, Alejandra Molina, Aaron Donsbach, Michael Terry, and Carrie J Cai. 2022. Promptmaker: Prompt-based prototyping with large language models. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI). 1–8. https://doi.org/10.1145/3491101.3503564
@@
- Jiang et al. (2023) Peiling Jiang, Jude Rayan, Steven P Dow, and Haijun Xia. 2023. Graphologue: Exploring large language model responses with interactive diagrams. In ACM Symposium on User Interface Software and Technology (UIST). 1–20. https://doi.org/10.1145/3586183.3606737
@@
- Jiang et al. (2020) Zhengbao Jiang, Frank F Xu, Jun Araki, and Graham Neubig. 2020. How can we know what language models know? Transactions of the Association for Computational Linguistics 8 (2020), 423–438. https://doi.org/10.1162/tacl_a_00324
@@
- Joshi et al. (2020) Pratik Joshi, Sebastin Santy, Amar Budhiraja, Kalika Bali, and Monojit Choudhury. 2020. The state and fate of linguistic diversity and inclusion in the NLP world. In Annual Meeting of the Association for Computational Linguistics (ACL). 6282–6293. https://doi.org/10.18653/v1/2020.acl-main.560
@@
- Kery et al. (2017) Mary Beth Kery, Amber Horvath, and Brad A Myers. 2017. Variolite: Supporting exploratory programming by data scientists.. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI), Vol. 10. 3025453–3025626. https://doi.org/10.1145/3025453.3025626
@@
- Kery and Myers (2017) Mary Beth Kery and Brad A Myers. 2017. Exploring exploratory programming. In IEEE Symposium on Visual Languages and Human-Centric Computing (VL/HCC). IEEE, 25–29. https://doi.org/10.1109/VLHCC.2017.8103446
@@
- Kim et al. (2016) Miryung Kim, Thomas Zimmermann, Robert DeLine, and Andrew Begel. 2016. The emerging role of data scientists on software development teams. In International Conference on Software Engineering (ICSE). 96–107. https://doi.org/10.1145/2884781.2884783
@@
- Ko et al. (2011) Amy J Ko, Robin Abraham, Laura Beckwith, Alan Blackwell, Margaret Burnett, Martin Erwig, Chris Scaffidi, Joseph Lawrance, Henry Lieberman, Brad Myers, et al. 2011. The state of the art in end-user software engineering. ACM Computing Surveys (CSUR) 43, 3 (2011), 1–44. https://doi.org/10.1145/1922649.1922658
@@
- Ko et al. (2015) Amy J Ko, Thomas D LaToza, and Margaret M Burnett. 2015. A practical guide to controlled experiments of software engineering tools with human participants. Empirical Software Engineering 20 (2015), 110–141. https://doi.org/10.1007/s10664-013-9279-3
@@
- Landis and Koch (1977) J Richard Landis and Gary G Koch. 1977. The measurement of observer agreement for categorical data. biometrics (1977), 159–174. https://doi.org/10.2307/2529310
@@
- LaToza et al. (2006) Thomas D LaToza, Gina Venolia, and Robert DeLine. 2006. Maintaining mental models: a study of developer work habits. In International Conference on Software Engineering (ICSE). 492–501. https://doi.org/10.1145/1134285.1134355
@@
- Liang et al. (2024) Jenny T Liang, Chenyang Yang, and Brad A Myers. 2024. A large-scale survey on the usability of AI programming assistants: Successes and challenges. In IEEE/ACM International Conference on Software Engineering (ICSE). 1–13. https://doi.org/10.1145/3597503.3608128
@@
- Liang et al. (2022) Jenny T Liang, Thomas Zimmermann, and Denae Ford. 2022. Understanding skills for OSS communities on GitHub. In ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE). 170–182. https://doi.org/10.1145/3540250.3549082
@@
- Lin (2024) Zhicheng Lin. 2024. How to write effective prompts for large language models. Nature Human Behaviour 8, 4 (2024), 611–615. https://doi.org/10.1038/s41562-024-01847-2

### v2

@@ 5.6. Evaluation
The participants had custom tasks that did not have a quantified notion of correctness, making the evaluation difficult (Parnin et al., 2023; Dolata et al., 2024). For building ML-enabled systems, many evaluations are quantitative metrics (Wan et al., 2019; Giray, 2021; Nahar et al., 2022), while prompt engineering uses manual inspection (Gero et al., 2022; Mishra et al., 2025). Participants performed manual tests by qualitatively assessing FM outputs (P2, P4, P6, P7, P8, P12, P13, P14, P16, P17, P18, P19, P20) as it was difficult to “quantify the concept of good” (P10). For example, P2, who created a voice memo summarization tool, did the following to test his application: “I would…talk to my phone for like 10 minutes just to fine-tune and see, ‘Does the summary make sense?”’ (P2).
@@
Manual testing was difficult to scale and quantify progress (P6, P15): “It is extremely hard to realize the impact of the [changes] without…a person who is very finely checking out the differences” (P15), especially in the face of regressions. Participants supplemented this with quantitative metrics via scaled programmatic testing (P1, P5, P8, P9, P11, P13, P14, P16, P18): “I run [my prompt] manually on 3 to 4 examples… [When] I have a proper evaluation setup, I run on 50 to 100 cases. Those 50 cases make me more confident” (P18). This involved using other FMs to generate scores for qualitative constructs (P4, P7, P9, P10, P15, P19), like safety (Furmakiewicz et al., 2024): “I have a scoring mechanism with…an evaluation agent [where the] score is from 0 to 100” (P19). However, FM-based evaluations could also introduce errors: “[The evaluation agent] itself is an LLM…So 10% of the time, if the test case is failing, it’s because the evaluation agent wasn’t able to score it correctly” (P19).
[…]
@@
Participants tested their prompts at varying scopes. For applications with multiple prompts, participants tested individual components of the prompt program (P8, P16, P19, P20), akin to unit testing: “Usually, each test case that we had written was aimed at testing one specific part of the prompt” (P20). The participants also described performing integration tests (P2, P8, P9, P11, P12, P13, P14, P16, P18, P19): “We evaluated each of [the prompts] separately, and now we are evaluating that as a collection” (P11). However, this type of testing was difficult, as combining prompts could yield unexpected interactions.
6. Discussion
We discuss how our findings relate to prompt engineering and traditional software engineering (Section 6.1), their implications (Section 6.2), and takeaways for various stakeholders (Section 6.3).
6.1. Prompt Programming: Where Software Engineering Meets Prompt Engineering
[…]
Implications for evolving FM capabilities
[…]
@@ 6.2. Implications
Developing ML-enabled systems requires a variety of expertise such as software engineering, data science, and math (Wan et al., 2019; Giray, 2021). This expertise is captured across team members, which can introduce collaboration challenges (Nahar et al., 2022). Though prompt programming involves both ML and programming knowledge, this expertise is now centralized in a single individual. This could reduce friction between different roles. However, in addition to traditional software engineering expertise like debugging, testing, or knowledge in specific programming languages (Liang et al., 2022; Baltes and Diehl, 2018), prompt programmers must develop new skills as they engage in novel activities like building datasets and debugging model errors with data (Qian et al., 2024).
Since prompt programmers are essentially ML practitioners, these programmers encounter many of the long-standing challenges contended by the ML and NLP communities, like model drift (Chen et al., 2024), language and cultural biases in datasets and models (Joshi et al., 2020; Santy* et al., 2023; Pei and Jurgens, 2023), and annotator subjectivity (Aroyo and Welty, 2015). Close collaboration between the software engineering and ML communities is paramount to inform techniques that address these challenges in prompt programming.
@@
The rapid iteration involved in prompt programming means that prompt programmers must make sense of a deluge of information. For each iteration, the participant must track the prompt, the change being made, the FM’s output, and associated metrics. However, participants reported having difficulty making sense of all of this information (P3, P6, P18, P19). This parallels exploratory programming, which is also highly iterative and unsystematic in nature (Kery and Myers, 2017) and requires tracking of many artifacts, including source code, input data, output data, code snippets, and analysis progress (Kery et al., 2017).
Are the “best practices” best practices?
In our study, participants’ mental models on FMs conflicted, such as whether few-shot learning (Brown et al., 2020) was beneficial for prompt performance. Although the inclusion of examples is listed within OpenAI’s official prompting guide (ope, 2024c), some of our participants (P7, P13, P19) were skeptical of this practice after observing the FM. This underscores the importance of the observations of developers in prompt programming, especially in the face of evolving model behaviors (Chen et al., 2024; Ma et al., 2024). However, such knowledge and observations are rarely shared, are siloed in informal social media posts (Parnin et al., 2023), and are not systematically collected and centralized.
6.3. Takeaways
For practitioners and educators
The lack of standardized best practices and unreliable mental models of prompt programmers pose challenges in teaching and learning prompt programming. Future work should investigate methods to effectively teach prompt programming. In the meantime, educators should consider the following recommendations, which are summarized in Figure 5.
[…]
@@ 6.3. Takeaways
“[Prompt programming] will require any…developer to adjust their mindset… It is much more scientific and experimental than traditional software development” (P15). Participants struggled at each step, from requirements to evaluation, as existing developer tools were not designed to handle the challenges of prompt programming. We follow Hassan et al. (2024) and Parnin et al. (2023) in underscoring the importance of developer tools for prompt programming. Such explorations have begun in earnest (e.g., Mishra et al., 2025; Wu et al., 2022a; Jiang et al., 2022; Arawjo et al., 2023). Future tool ecosystems should support the full range of activities in prompt programming, from interfaces for rapid prototyping, dataset curation, and data sensemaking, to regression testing, prompt chaining, collaboration, and more. We elaborate on potential prompt programming tool directions below.
Given the large volume of information prompt programmers process per iteration (e.g., examples, prompt modifications, and FM output changes), future work should explore improved methods for versioning and navigating prompt iterations and artifacts. This could provide rich information on prompt provenance and could be used for automatic prompt program repair. A key challenge is deciding what to version and how to represent both prompt and FM output changes. Because prompt changes are made with varying granularity, from editing single words to changing prompting strategies, research is needed to understand how to visualize these changes between versions. Meanwhile, research should study new visualizations for FM output changes. This is challenging as prompt programmers qualitatively compare the same prompt across multiple examples (Gero et al., 2022).
Future work is also needed for the reuse of prompts, since the fragility of prompts and the unique qualities of each FM make it difficult to reuse prompts without modification. Reuse at higher levels of abstraction (e.g., template-based approaches (White et al., 2023, 2024)), and community-based solutions for prompt sharing (e.g., ShareGPT (sha, 2024), LangSmith Hub (lan, 2024), Wordflow (Wang et al., 2024)) could be promising avenues for prompt reuse. Research is needed to understand how to retrieve relevant prompts. Prompt programmers could query prompts based on different information, ranging from specific keywords in the input and FM output to general FM behavior or performance metrics. Another challenge is ensuring that the recommendations can generalize across diverse, situation-specific contexts of different prompt programmers (e.g., identifying optimal wording for specific applications).
7. Conclusion
Foundation models (FMs) enable programmers to write natural language prompts that power AI-driven software features. We study this phenomenon—prompt programming—via 20 interviews with developers across various contexts, using Straussian grounded theory to analyze the process. We identified 15 observations about prompt programming barriers. We find that prompts are the product of the developer’s understanding of FM and the developer’s observations of the FM’s behavior on the task. We also find strong parallels between the prompt programming process and the traditional software engineering process, but also significant differences. Our findings inform both researchers developing prompt programming tools and practitioners who use them.
@@ 7. Conclusion
To facilitate replication, our supplemental materials are available on FigShare (Liang et al., 2025),including our survey instrument, interview protocol, codebook, and the 33 reviewed papers.
@@ Acknowledgements.
We thank our participants for their insights. We express gratitude to Chenyang Yang, Nadia Nahar, Travis Breaux, and Christian Kaestner for their feedback. Last but not least, we give a special thanks to MeiReferences
@@
- Arawjo et al. (2023) Ian Arawjo, Priyan Vaithilingam, Martin Wattenberg, and Elena Glassman. 2023. ChainForge: An open-source visual programming environment for prompt engineering. In ACM Symposium on User Interface Software and Technology (UIST). 1–3. doi:10.1145/3586182.3616660
- Aroyo and Welty (2015) Lora Aroyo and Chris Welty. 2015. Truth is a lie: Crowd truth and the seven myths of human annotation. AI Magazine 36, 1 (2015), 15–24. doi:10.1609/aimag.v36i1.2564
- Arpteg et al. (2018) Anders Arpteg, Björn Brinne, Luka Crnkovic-Friis, and Jan Bosch. 2018. Software engineering challenges of deep learning. In Euromicro Conference on Software Engineering and Advanced Applications (SEAA). IEEE, 50–59. doi:10.1109/SEAA.2018.00018
- Baltes and Diehl (2018) Sebastian Baltes and Stephan Diehl. 2018. Towards a theory of software development expertise. In ACM Joint Meeting on European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE). 187–200. doi:10.1145/3236024.3236061
- Bender et al. (2021) Emily M Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell. 2021. On the dangers of stochastic parrots: Can language models be too big?. In ACM conference on Fairness, Accountability, and Transparency (FAccT). 610–623. doi:10.1145/3442188.3445922
- Betker et al. (2023) James Betker, Gabriel Goh, Li Jing, Tim Brooks, Jianfeng Wang, Linjie Li, Long Ouyang, Juntang Zhuang, Joyce Lee, Yufei Guo, et al. 2023. Improving image generation with better captions. 2, 3 (2023), 8. doi:papers/dall-e-3. pdf
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Advances in Neural Information Processing Systems (NeurIPS) 33 (2020), 1877–1901.
@@
- Chavan et al. (2024) Omkar Sandip Chavan, Divya Dilip Hinge, Soham Sanjay Deo, Yaxuan Wang, and Mohamed Wiem Mkaouer. 2024. Analyzing developer-ChatGPT conversations for software refactoring: An exploratory study. In International Conference on Mining Software Repositories (MSR). 207–211. doi:10.1145/3643991.3645082
- Chen et al. (2024) Lingjiao Chen, Matei Zaharia, and James Zou. 2024. How is ChatGPT’s behavior changing over time? Harvard Data Science Review 6, 2 (2024). doi:10.1162/99608f92.5317da47
@@
- Cole and Gillies (2022) Tom Cole and Marco Gillies. 2022. More than a bit of coding:(un-) Grounded (non-) theory in HCI. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI). 1–11. doi:10.1145/3491101.3516392
- Corbin and Strauss (2015) Juliet Corbin and Anselm Strauss. 2015. Basics of qualitative research. Vol. 14. Sage. doi:10.1177/1094428108324514
- Dijkstra (1979) Edsger W Dijkstra. 1979. On the foolishness of ”natural language programming”. (1979), 51–53. doi:10.1007/BFb0014656
- Dijkstra (1982) Edsger W Dijkstra. 1982. On the role of scientific thought. Selected writings on computing: A personal perspective (1982), 60–66. doi:10.1007/978-1-4612-5695-3_12
- Dolata and Crowston (2023) Mateusz Dolata and Kevin Crowston. 2023. Making sense of AI systems development. IEEE Transactions on Software Engineering (TSE) (2023). doi:10.1109/TSE.2023.3338857
- Dolata et al. (2024) Mateusz Dolata, Norbert Lange, and Gerhard Schwabe. 2024. Development in times of hype: How freelancers explore generative AI?. In IEEE/ACM International Conference on Software Engineering (ICSE). 1–13. doi:10.1145/3597503.3639111
@@
- Giray (2021) Görkem Giray. 2021. A software engineering perspective on engineering machine learning systems: State of the art and challenges. Journal of Systems and Software 180 (2021), 111031. doi:10.1016/j.jss.2021.111031
- Glaser and Strauss (2017) Barney Glaser and Anselm Strauss. 2017. Discovery of grounded theory: Strategies for qualitative research. Routledge. doi:10.4324/9780203793206
@@
- Hassan et al. (2024) Ahmed E Hassan, Dayi Lin, Gopi Krishnan Rajbahadur, Keheliya Gallaba, Filipe Roseiro Cogo, Boyuan Chen, Haoxiang Zhang, Kishanthan Thangarajah, Gustavo Oliva, Jiahuei Lin, et al. 2024. Rethinking software engineering in the era of foundation models: A curated catalogue of challenges in the development of trustworthy FMware. In Companion Proceedings ACM International Conference on the Foundations of Software Engineering. 294–305. doi:10.1145/3663529.3663849
@@
- Jacovi and Goldberg (2020) Alon Jacovi and Yoav Goldberg. 2020. Towards faithfully interpretable NLP systems: How should we define and evaluate faithfulness?. In Annual Meeting of the Association for Computational Linguistics (ACL), Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel Tetreault (Eds.). 4198–4205. doi:10.18653/v1/2020.acl-main.386
@@
- Jiang et al. (2022) Ellen Jiang, Kristen Olson, Edwin Toh, Alejandra Molina, Aaron Donsbach, Michael Terry, and Carrie J Cai. 2022. Promptmaker: Prompt-based prototyping with large language models. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI). 1–8. doi:10.1145/3491101.3503564
@@
- Jiang et al. (2023) Peiling Jiang, Jude Rayan, Steven P Dow, and Haijun Xia. 2023. Graphologue: Exploring large language model responses with interactive diagrams. In ACM Symposium on User Interface Software and Technology (UIST). 1–20. doi:10.1145/3586183.3606737
@@
- Jiang et al. (2020) Zhengbao Jiang, Frank F Xu, Jun Araki, and Graham Neubig. 2020. How can we know what language models know? Transactions of the Association for Computational Linguistics 8 (2020), 423–438. doi:10.1162/tacl_a_00324
@@
- Joshi et al. (2020) Pratik Joshi, Sebastin Santy, Amar Budhiraja, Kalika Bali, and Monojit Choudhury. 2020. The state and fate of linguistic diversity and inclusion in the NLP world. In Annual Meeting of the Association for Computational Linguistics (ACL). 6282–6293. doi:10.18653/v1/2020.acl-main.560
@@
- Kery et al. (2017) Mary Beth Kery, Amber Horvath, and Brad A Myers. 2017. Variolite: Supporting exploratory programming by data scientists.. In ACM CHI Conference on Human Factors in Computing Systems Extended Abstracts (CHI), Vol. 10. 3025453–3025626. doi:10.1145/3025453.3025626
@@
- Kery and Myers (2017) Mary Beth Kery and Brad A Myers. 2017. Exploring exploratory programming. In IEEE Symposium on Visual Languages and Human-Centric Computing (VL/HCC). IEEE, 25–29. doi:10.1109/VLHCC.2017.8103446
@@
- Kim et al. (2016) Miryung Kim, Thomas Zimmermann, Robert DeLine, and Andrew Begel. 2016. The emerging role of data scientists on software development teams. In International Conference on Software Engineering (ICSE). 96–107. doi:10.1145/2884781.2884783
@@
- Ko et al. (2011) Amy J Ko, Robin Abraham, Laura Beckwith, Alan Blackwell, Margaret Burnett, Martin Erwig, Chris Scaffidi, Joseph Lawrance, Henry Lieberman, Brad Myers, et al. 2011. The state of the art in end-user software engineering. ACM Computing Surveys (CSUR) 43, 3 (2011), 1–44. doi:10.1145/1922649.1922658
@@
- Ko et al. (2015) Amy J Ko, Thomas D LaToza, and Margaret M Burnett. 2015. A practical guide to controlled experiments of software engineering tools with human participants. Empirical Software Engineering 20 (2015), 110–141. doi:10.1007/s10664-013-9279-3
@@
- Landis and Koch (1977) J Richard Landis and Gary G Koch. 1977. The measurement of observer agreement for categorical data. biometrics (1977), 159–174. doi:10.2307/2529310
@@
- LaToza et al. (2006) Thomas D LaToza, Gina Venolia, and Robert DeLine. 2006. Maintaining mental models: a study of developer work habits. In International Conference on Software Engineering (ICSE). 492–501. doi:10.1145/1134285.1134355
- Liang et al. (2025) Jenny T Liang, Melissa Lin, Nikitha Rao, and Brad Myers. 2025. Supplemental Materials to ”Prompts are programs too! Understanding how developers build software containing prompts”. doi:10.6084/m9.figshare.24210468
@@
- Liang et al. (2024) Jenny T Liang, Chenyang Yang, and Brad A Myers. 2024. A large-scale survey on the usability of AI programming assistants: Successes and challenges. In IEEE/ACM International Conference on Software Engineering (ICSE). 1–13. doi:10.1145/3597503.3608128
@@
- Liang et al. (2022) Jenny T Liang, Thomas Zimmermann, and Denae Ford. 2022. Understanding skills for OSS communities on GitHub. In ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE). 170–182. doi:10.1145/3540250.3549082
@@
- Lin (2024) Zhicheng Lin. 2024. How to write effective prompts for large language models. Nature Human Behaviour 8, 4 (2024), 611–615. doi:10.1038/s41562-024-01847-2